Lehighton Chalk Company manufactures sidewalk chalk, which it sells online by the box at $25 per unit. Lehighton uses an actual costing system, which means that the actual costs of direct material, direct labor, and manufacturing overhead are entered into work-in-process inventory. The actual application rate for manufacturing overhead is computed each year; actual manufacturing overhead is divided by actual production (in units) to compute the application rate. Information for Lehighton’s first two years of operation is as follows:

Year 1 Year 2
Sales (in units) 2,600 2,600
Production (in units) 3,100 2,100
Production costs:
Variable manufacturing costs $ 15,500 $ 10,500
Fixed manufacturing overhead 18,600 18,600
Selling and administrative costs:
Variable 10,400 10,400
Fixed 9,400 9,400

Selected information from Lehighton’s year-end balance sheets for its first two years of operation is as follows:

LEHIGHTON CHALK COMPANY
Selected Balance Sheet Information
Based on absorption costing End of Year 1 End of Year 2
Finished-goods inventory $ 5,500 $ 0
Retained earnings 11,100 19,000
Based on variable costing End of Year 1 End of Year 2
Finished-goods inventory $ 2,500 $ 0
Retained earnings 8,100 19,000

Required:

  1. Reconcile Lehighton’s operating income reported under absorption and variable costing, during each year, by comparing the following two amounts on each income statement:

  • Cost of goods sold
  • Fixed cost (expensed as a period expense)

Year 1 Year 2 Subtotal Total 0 $ 0 Difference in operating income

  1. What was Lehighton’s total operating income across both years under absorption costing and under variable costing?

Total Operating Income Absorption costing Variable costing

  1. What was the total sales revenue across both years under absorption costing and under variable costing?

Total Sales Revenue Absorption costing Variable costing

  1. What was the total of all costs expensed on the operating income statements across both years under absorption costing and under variable costing?

Costs Expensed Absorption costing Variable costing

  1. Subtract the total costs expensed across both years [requirement (4)] from the total sales revenue across both years [requirement (3)]: (a) under absorption costing and (b) under variable costing.

Amount Absorption costing Variable costing

  1. Considering the results obtained in requirements 1-5 above, select which of the following statements (is) are true by selecting an "X".

Year 1 Year 2 Subtotal Total 0 $ 0 Difference in operating income
Total Operating Income Absorption costing Variable costing
Total Sales Revenue Absorption costing Variable costing
Costs Expensed Absorption costing Variable costing
Amount Absorption costing Variable costing
Sales revenue is different depending on the costing method used. Timing is the key in distinguishing between absorption and variable costing. Since Lehighton's combined operating income, across the two-year period, is the same under both absorption and variable costing, then the operating income must be the same within each year under both methods. The difference between absorption and varible costing is caused by the timing with which expenses are recognized.
Answer #1

Solution

Lehighton Chalk Company                  

  1. Reconcile

Reconciliation:

Year 1

Year 2

Cost of goods sold under absorption costing

$28,600

$34,600

Less: Variable manufacturing costs under variable costing

$13,000

$13,000

sub total

$15,600

$21,600

Less: Fixed manufacturing overhead as period expense under variable costing

$18,600

$18,600

Total

($3,000)

$3,000

Operating income under variable costing

$13,600

$13,600

Less: operating income under absorption costing

$16,600

$10,600

Difference in operating income

($3,000)

$3,000

  1. Computation of operating income across both years under absorption costing and under variable costing:

Operating Income

Year 1

Absorption Costing

27,200

Variable Costing

$23,380

Computations:

Absorption Costing Income Statement

Year 1

Year 2

Sales (2,600 units)

65,000

Sales (2,600 units)

$65,000

Less: Cost of Goods Sold

Less: Cost of Goods Sold

Beginning inventory

$0

Beginning inventory

(500 x $11)

$5,500

Add: cost of goods manufactured

(3,100 x $8.50)

$34,100

Add: cost of goods manufactured

(1,900 x 13.86)

$29,100

Goods Available for Sale

$34,100

Goods Available for Sale

$34,600

Less: ending inventory

(500 x $11)

$5,500

Less: ending inventory

0

$0

Cost of goods sold

$28,600

Cost of goods sold

$34,600

Gross Margin

$36,400

Gross Margin

$30,400

Less: selling and administrative expenses

Less: selling and administrative expenses

Variable selling expenses

$10,400

Variable selling expenses

$10,400

Fixed Selling expenses

$9,400

Fixed Selling expenses

$9,400

Total selling and administrative expenses

$19,800

Total selling and administrative expenses

$19,800

Operating Income

$16,600


Operating Income

$10,600

Computations:

Cost of goods manufactured, Year 1 –

Variable + fixed overhead

Variable overhead = 15,500/3,100 = $5

Fixed overhead = 18,600/3,100 = $6

Total cost of goods manufactured per unit = $11

Cost of goods manufactured, Year 2 –

Variable = $5

Fixed = 18,600/2,100 units = $8.86

Total cost of goods manufactured per unit = $13.86

  1. Operating income statements for both years using variable costing method:

Variable Costing Income Statement

Year 1

Year 2

sales

$65,000

sales

$65,000

Variable expenses:

Variable expenses:

Variable cost of goods manufactured

Variable cost of goods manufactured

Beg. Inventory

0

Beg. Inventory

$2,500

Variable cost of goods manufactured

$15,500

Variable cost of goods manufactured

$10,500

Less: ending inventory

$2,500

Less: ending inventory

0

Variable cost of goods sold

$13,000

Variable cost of goods sold

$13,000

Manufacturing margin

$52,000

Manufacturing margin

$52,000

Less: Variable selling expenses

$10,400

Less: Variable selling expenses

$10,400

Contribution margin

$41,600

Contribution margin

$41,600

Fixed costs:

Fixed costs:

Manufacturing overhead

$18,600

Manufacturing overhead

$18,600

Selling expenses

$9,400

Selling expenses

$9,400

Total fixed expenses

$28,000

Total fixed expenses

$28,000

Net Income

$13,600

Net Income

$13,600

  1. Total sales revenue across both years under absorption costing and under variable costing:

Total Sales Revenue

Absorption Costing

$130,000

Variable Costing

$130,000

  1. Total of all costs expensed on the operating income statements across both years under absorption costing and under variable costing:

Costs Expensed

Absorption Costing

$102,800

Variable Costing

$102,800

Computations:

Total cost expensed = total cost of goods sold for two years + selling and administrative expenses for two years

Absorption Costing –

= (28,600 + 34,600) + (19,800 + 19,800) = $102,800

Variable costing –

= (13,000 + $13,000) + (10,400 + 10,400) + (28,000 + 28,000) = $102,800

  1. Subtract total costs expensed under both years from the total sales revenue across both years under absorption costing and variable costing:

Amount

Absorption Costing

$27,200

Variable Costing

$27,200

  1. Considering the results -

Select which of the following statements is true by selecting an 'X':

Sales revenue is different depending on the costing method used

FALSE

Timing is the key in distinguishing between absorption and variable costing

X

Since Lehighton's combined operating income, across the two-year period is the same under both absorption and variable costing, then the operating income must be the same within each year under both the methods

FALSE

The difference between absorption and variable costing is caused by the timing with which expenses are recognized

X

Check the reconciliation:

Year

Change in inventory in units

Actual fixed overhead

Difference in fixed overhead

Absorption - variable costing operating income difference

1

500

+

$6

$3,000

$3,000

2

-500

-

$6

($3,000)

($3,000)